Solve for exact solutions over [0,2pie) invterval
sec(x/2)=cos(x/2)

Answer by Alan3354(69443)

Solve for exact solutions over [0,2pie) invterval
sec(x/2)=cos(x/2)
--------------
Use the Greek letter pi, not the dessert, pie.
--------
sec(x/2)=cos(x/2)
1/cos(x/2)=cos(x/2)
cos^2(x/2) = 1
cos(x/2) = -1,1
x/2 = 0, pi
x = 0, 2pi